OpenStudy (anonymous):

Which equation is equivalent to 2(3x - 1) = 10 + 2x?

OpenStudy (anonymous):

what are the other equations?

OpenStudy (anonymous):

this is the equation : 2(3x-1)=10+2x

OpenStudy (anonymous):

The first thing you need to do is to distribute the 2: 6x-2=10+2x Next, subtract 2x from both sides to collect all terms with x on one side of the equation: 4x-2=10 This equation matches with answer b. So, b is your answer.

OpenStudy (anonymous):

It says "which" equation implying that there's others.

OpenStudy (anonymous):

that is if ur answers are a. 8x - 2 = 10 b. 4x - 2 = 10 c. 6x - 11 + 2x d. 6x - 1 = 10 + 2X i did this a while ago